# Quantum-Inspired Weight-Constrained Neural Network: Reducing Variable Numbers by 100x Compared to Standard Neural Networks

### v1

@@ Abstract
Although quantum machine learning has shown great promise, the practical application of quantum computers remains constrained in the noisy intermediate-scale quantum era. To take advantage of quantum machine learning, we investigate the underlying mathematical principles of these quantum models and adapt them to classical machine learning frameworks. Specifically, we develop a classical weight-constrained neural network that generates weights based on quantum-inspired insights. We find that this approach can reduce the number of variables in a classical neural network by a factor of 135 while preserving its learnability. In addition, we develop a dropout method to enhance the robustness of quantum machine learning models, which are highly susceptible to adversarial attacks. This technique can also be applied to improve the adversarial resilience of the classical weight-constrained neural network, which is essential for industry applications, such as self-driving vehicles. Our work offers a novel approach to reduce the complexity of large classical neural networks, addressing a critical challenge in machine learning.
Introduction
@@
Emerging quantum machine learning techniques offer promising solutions to some challenges posed by large AI models [13, 14, 15, 16, 17, 18]. The advantage of quantum speed-ups has been demonstrated in various machine learning tasks [14, 19], including finding concise functions to fit exponentially large datasets [20], learning a particular data distribution [18], discovering unknown kernel functions [21], and extracting insights from experimental data points using exponentially fewer data compared to classical approaches [22]. These advances can significantly reduce the computational time required to train AI models. Beyond speed, quantum machine learning offers memory advantages by taking advantage of the exponential storage capability of quantum computers. For example, quantum neural networks with amplitude encoding require far fewer variables than their classical counterparts [23]. Consequently, the vast number of parameters in large AI models can be effectively reduced by employing quantum or hybrid quantum-classical models.
@@
Driven by these advantages, quantum machine learning (QML) has become an active area of research [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Although substantial progress has been made in the development of QML algorithms, there are still few real-world applications that leverage real quantum computers [35, 36, 37, 38, 39]. This gap largely results from the quantum noise issue, which prevents reliable results when many qubits are used [40, 41]. Addressing this noise challenge is complex, and achieving a large-scale, noise-free quantum computer is expected to take considerable time. In light of these obstacles, the development of quantum-inspired classical algorithms that address the challenges of classical machine learning is a valuable direction to advance machine learning techniques [42, 43, 44, 45, 46, 47, 48].
[…]
@@
In this work, we present a method to reduce the number of variables in classical deep-learning neural networks by drawing inspiration from quantum neural networks that employ angle and amplitude encodings. Our analysis reveals that using a parameterized quantum circuit with angle encoding is mathematically equivalent to constructing a polynomial function, where each term is expressed as a continuous product of the cosine series components [47]. In contrast, a quantum circuit with amplitude encoding is designed to establish correlations between elements in the input data. We note that the number of variables in the correlation matrix is determined by the number of parameterized gates rather than the dimensionality of the input data, implying that the quantum circuit with amplitude encoding needs fewer variables compared to the classical neural network. These insights motivate us to develop an algorithm capable of generating thousands of weights using only dozens of variables. Specifically, we utilize a combinatorial approach to create numerous sets of angle combinations. Each combination is then used to construct weights through the continuous product of trigonometric functions. We validate this weight-constrained method by examining its performance on four datasets: MNIST, Fashion MNIST, CIFAR, and traffic datasets. Our results demonstrate that the weight-constrained method maintains the learnability of the classical neural network while substantially reducing the number of variables. This approach holds promise for scaling down variable numbers in large AI models such as GPT and BERT. Furthermore, we develop a novel method to enhance the adversarial resilience of the quantum neural network and the classical weight-constrained neural network models by randomly dropping parameterized angles during prediction. We demonstrate that such a method can significantly increase resilience against adversarial attacks.
Quantum neural networks
@@
To develop quantum-inspired classical neural networks, we explore the underlying mathematics of quantum neural networks with angle and amplitude encodings [1]. In this work, we elucidate how these two types of quantum neural networks select features. After building on this mathematical foundation, we develop a quantum-inspired classical neural network, which utilizes 100 times fewer variables than a standard classical neural network.
@@
We start to analyze the quantum neural network with angle encoding, which adopts the input value as rotation angles in the quantum gates, as shown in Fig. 1a. The nonlinear relationship between the output of a quantum circuit and the input data ${\bf z}$ is constructed by applying different rotation gates and CNOT gates in this circuit. Without loss of generality, we consider a quantum circuit shown in Fig. 1a, which has one data-upload layer $\tilde{R}_{0}$, $M$ rotation layers, and $M$ CNOT layers. The rotation operator $R_{\gamma}$ in Fig. 1a is chosen from the group $\{R_{x},R_{y},R_{z}\}$. Then, the final quantum state of this quantum circuit is
@@
where $\tilde{R}_{i}$ denotes a matrix constructed from the tensor product of different single-gate rotation matrices within one layer, and $\tilde{C}_{i}$ denotes a matrix constructed from the tensor product of two-gate CNOT matrices. If we adopt measurements $A$ with Pauli operators, it is straightforward to find that the output $O$ of a quantum circuit is a polynomial function, which is expressed as
$O$ $=$ $\langle\psi|A|\psi\rangle=\sum_{s}f_{s}({\bf\theta},{\bf z}),$ (2)
$f_{s}({\bf\theta},{\bf z})$ $=$ $c_{s}\prod_{i}\mathrm{cos}(\theta_{i}/2+q_{i})\prod_{j}\mathrm{cos}(\theta_{j}/2+q_{j})\times$ (3)
$\prod_{n}\mathrm{cos}(z_{n}/2+q_{n})\prod_{m}\mathrm{cos}(z_{m}/2+q_{m}),$
where $c_{s}$ belongs to {$1,-1$}. The symbol $q_{i}$ denotes 00 or $\pi/2$. Equation (2) shows that the quantum neural network with angle encoding creates a complicated activation function, which could select features better than the classical activation functions, such as ReLU and Tanh. We refer the reader to our previous work [47] for a deeper understanding of this quantum activation function.
Quantum neural network with amplitude encoding
@@
where $|b_{i}\rangle$ denotes a pure quantum state, i.e., $|0000\rangle$ and $|0001\rangle$. $|{\bf z}|$ is the norm value of the input vector ${\bf z}$. This initial quantum state can be transformed to other states by applying rotation and CNOT gates. For simplicity, we can use an operator $U$ to represent these operations as shown in Fig. 1b. Then, the final quantum state is written as
@@
We can eliminate the factor $\frac{1}{|{\bf z}|^{2}}$ because it is equivalent to data normalization. A key observation from Eq. (7) is that the quantum circuit creates a correlation matrix between any two elements in the input data. For example, $W_{1,2}$ represents the correlation between the first and second input features, and $W_{1,100}$ denotes the correlation between the first and the one hundredth input features. If all elements in $W$ are nonzero, the construction of features incorporates long-range and short-range correlations[21]. It is important to emphasize that the quantum convolutional neural network proposed in Ref. 49, which applies convolutional operations between neighboring qubits, inherently creates long-range entanglement among real-world input features. This characteristic fundamentally differentiates the feature selection in the quantum convolutional neural network from that in the classical convolutional neural network, as the latter typically relies on spatially localized filters.
Figure 1c shows a hybrid quantum-classical neural network (HNN) that we adopt to perform the task, classifying whether a data point is above or below a boundary. In this setup, the quantum circuit is responsible for feature constructions, while a fully connected classical neural network with a linear activation function performs the classification. The quantum circuits for angle encoding and amplitude encoding are depicted in Figs. 1d and 1e, respectively. The two-qubit operators $U$ and $V$ are sketched in supplementary note 1. As shown in Figs. 1f-1i, we evaluate the performance of the hybrid network on two different boundaries: $y_{1}=x/2$ and $y_{2}=x(x-0.5)(x+0.5)/2$, using angle and amplitude encodings. The black curve denotes the boundary, the green dot denotes the predicted value above the boundary, and the red dot denotes the predicted value below the boundary. The QNN with amplitude encoding struggles to identify both boundaries because the features derived from terms like $x^{2}$, $xy$, and $y^{2}$ are insufficient to characterize $y_{1}$ and $y_{2}$. However, the QNN with angle encoding successfully identifies the boundary for $y_{2}$ due to its ability to capture the third-order term (i.e., $x^{3}$) inherent in the problem. The higher-order term can be obtained from the Taylor expansion of the cosine function. The subpar performance of angle encoding for $y_{1}$ can be attributed to the excessive nonlinearity imposed by the QNN.
Although QNNs with amplitude encoding exhibit lower learnability compared to those with angle encoding, they offer a key advantage: the ability to represent large datasets using relatively few qubits, thereby requiring fewer variables. This benefit is reflected in the substantially reduced number of variables in the weight matrix $W$ shown in Eq. (7). In other words, the weight elements in $W$ are not independent, and the variable number of the weight matrix is determined by the number of rotation angles rather than the dimensionality of the matrix. This observation suggests that independent weights $W_{ij}$ may not be necessary for classical deep-learning neural networks.
Figure 2: The expressibility and trainability of a neural network using the weight-constrained method. a The distribution $P(z)$ of $z$, defined as $z=\sum_{i}w_{i}x_{i}$, where $w$ denotes the weight, and $x$ denotes the input features. b The inverse of Kullback-Leibler divergence $D_{\text{KL}}^{-1}$ as a function of $N$, where $N$ denotes the number of variables. The right axis shows the inverse of the standard derivation $\Delta^{-1}$. c The inverse of Kullback-Leibler divergence $D_{\text{KL}}^{-1}$ as a function of $r$, where $r$ denotes the number of variables that are selected from $N$ variables. The blue dashed line represents a fitted curve.
Quantum-inspired classical network: weight-constrained neural network
By exploring quantum neural networks with angle and amplitude encodings, we find the underlying mathematical principles of these two quantum models. Our key finding reveals that the weight elements in quantum neural networks are determined by parametric rotation angles through trigonometric functions. Building on this insight, we develop a classical weight-constrained neural network.
@@
In the classical weight-constrained neural network, we construct weight variables as follows. To generate $K$ different weights, we first define $N$ different variables, denoted as $\theta$. Then a combination process is performed, where $r$ variables are chosen from these $N$ variables, resulting in $C(N,r)$ possible combinations. Labeling the $k$-th combination as ${\bf\theta}^{(k)}$, the $k$-th weight $w_{k}$ is defined as
@@
When the weight-constrained method is applied to a classical neural network, it reduces the learnability of the network, which refers to the ability to learn a target function. To investigate the impact of this constraint on learnability, we examine two distinct two-layer neural networks. In the first classical network, 1000 input features, $x$, are connected to an output neuron, $y$, as follows:
@@
where $w_{i}$ represents the weight, sampled from the uniform distribution $U(-1,1)$. Here, $f(z)$ denotes a standard classical activation function, such as Tanh, Sigmoid, or ReLU. The second classical neural network has the same architecture as the first. However, in the second network, weights are determined using the weight-constrained method described in Eq. (8). Notably, the uniformly distributed bias term affects the distribution of $z$ in the same way for both networks and is therefore omitted from this analysis.
Figure 2a displays the distribution $P(z)$ of $z$ for both classical neural networks. The input features $x$ are sampled from a uniform distribution $U(-1,1)$. The gray shading represents the distribution for the classical neural network with weights sampled from $w\in U(-1,1)$. The colored curves correspond to the classical weight-constrained neural network for various values of $N$ and $r$, which are parameters used to construct the combination sets. As shown in Fig. 2a, both networks exhibit dome-shaped distributions centered at $z=0$. Including a uniformly distributed bias term does not affect the central position of these distributions. It is observed that the weight-constrained network displays a narrower distribution width and a more pronounced peak compared to the network with uniformly distributed weights. This result suggests that the weight-constrained method limits access to larger $z$ values, thereby reducing the learnability of the network. When the activation function is applied to $z$, the weight-constrained method impacts the expressibility of different activation functions differently. For example, the weight-constrained method has a smaller effect on the Tanh function compared to the ReLU function because larger values of $|z|$ are less essential for the Tanh function. Further details on this observation can be found in Supplementary Note 3.
@@
Figure 3: Training results of different neural networks. a The architecture of the hybrid quantum-classical neuron network (HNN) and the training results on the MNIST and Fashion MNIST (FMNIST) datasets. Here, ACC denotes accuracy. b The fully connected neuron network (FNN) architecture and the training results on the MNIST and FMNIST datasets. c The weight-constrained FNN architecture and the training results on the MNIST and FMNIST datasets. d The convolutional neural network (CNN) architecture and the training results on the MNIST and FMNIST datasets. e The architecture of the weight-constrained CNN and the training results on the MNIST and FMNIST datasets. f A summary of the ACC of these five models for MNIST and FMNIST datasets.
To quantify the reduction in learnability caused by the weight-constrained method, we employ the Kullback–Leibler divergence ($D_{\text{KL}}$), defined as
$D_{\text{KL}}=\sum_{z}p(z)\log\left(\frac{p(z)}{q(z)}\right),$ (13)
where $p(z)$ represents the distribution of $z$ for the network using the weight-constrained method, and $q(z)$ represents the distribution for the network with uniformly distributed weights. The value of $D_{\text{KL}}$ is inversely proportional to the expressibility of the network; thus, a higher expressibility corresponds to a larger $D^{-1}_{\text{KL}}$ [50].
Figures 2b and 2c depict the inverse of the Kullback–Leibler divergence ($D^{-1}_{\text{KL}}$, blue circles) as a function of $N$ (at $r=4$) and $r$ (at $N=20$), respectively. The result shows that the increase in $N$ or $r$ reduces the value of $D^{-1}_{\text{KL}}$, indicating a reduction in expressibility. Notably, the expressibility declines more rapidly with increasing $r$ compared to increasing $N$. Using a fitting approach, we find that $D^{-1}_{\text{KL}}$ decays as a function of $1/2^{r}$, which is shown as the blue dashed curve in Fig. 2c. This result highlights the importance of selecting a small value of $r$ to preserve strong expressibility.
For the Tanh and Sigmoid activation functions, the output $y$ exhibits a larger slope when $z$ is near zero, but the slope approaches zero for $|z|>2$. A larger slope generally indicates better trainability. To assess the trainability of the network using the weight-constrained method, we examine the standard deviation $\Delta$ of the distribution $p(z)$, represented by the red triangular symbols in Figures 2b and 2c. At $r=4$, $\Delta^{-1}$ is nearly constant with varying $N$, indicating that changes in $N$ do not significantly impact trainability. In contrast, $\Delta^{-1}$ shows a linear dependence but not an exponential dependence on $r$, implying that the phenomenon of quantum barren plateaus [51] is absent in the classical weight-constrained neural network. Moreover, the result in Fig. 2c indicates that the weight-constrained network with larger values of $r$ exhibits higher trainability. However, given the reduced expressibility at large $r$, it is crucial to optimize $r$ to balance expressibility and trainability. It is worth noting that these trainability analyses do not apply to the ReLU activation function because the slope for $z>0$ is independent of the value of $z$.
Practicality of the weight-constrained method
Having discussed the expressibility and trainability of classical neural networks employing the weight-constrained method, we now focus on its practical applications in classical fully connected and convolutional neural networks.
We begin by comparing the performance of the classical weight-constrained neural network with that of a standard classical fully connected neural network (FNN) and a hybrid neural network (HNN). The architectures of these three networks are depicted in Figs. 3a3c, respectively. All three networks consist of four layers, with 16 neurons in the first hidden layer and 64 neurons in the second hidden layer. Details on the quantum circuit design are provided in Supplementary Note 1. For the weight-constrained FNN, the weights in the first hidden layer are determined using Eq.(8). In addition to FNN, we evaluate the performance of the classical weight-constrained convolutional neural network (CNN) and its standard counterpart. Figure 3d illustrates a standard CNN architecture comprising one convolutional layer, one pooling layer, and two fully connected layers. The weight-constrained CNN, shown in Fig. 3e, shares the same architecture as the standard CNN, but the weight-constrained method is applied to the first fully connected layer.

### v2

@@ Abstract
Although quantum machine learning has shown great promise, the practical application of quantum computers remains constrained in the noisy intermediate-scale quantum era. To take advantage of quantum machine learning, we investigate the underlying mathematical principles of these quantum models and find that the quantum neural network with amplitude encoding is equivalent to a weight-constrained neural network. Motived by this discovery, we develop a classical weight-constrained neural network. We find that this approach can reduce the number of variables in a classical neural network by a factor of 135 while preserving its accuracy. In addition, we develop a dropout method to enhance the robustness of quantum machine learning models, which are highly susceptible to adversarial attacks. This technique can also be applied to improve the adversarial robustness of the classical weight-constrained neural network, which is essential for industry applications, such as self-driving vehicles. Our work offers a novel approach to reduce the complexity of large classical neural networks, addressing a critical challenge in machine learning.
I Introduction
@@
Emerging quantum machine learning techniques offer promising solutions to some challenges posed by large AI models [13, 14, 15, 16, 17, 18]. The advantage of quantum speed-ups has been demonstrated in various machine learning tasks [14, 19], including finding concise functions to fit exponentially large datasets [20], learning a particular data distribution [18], discovering unknown kernel functions [21], and extracting insights from experimental data points using exponentially fewer data compared to classical approaches [22]. These advances can significantly reduce the computational time required to train AI models. In addition to speed, quantum machine learning offers memory advantages by taking advantage of the exponential storage capability of quantum computers. For example, quantum neural networks with amplitude encoding require far fewer variables than their classical counterparts [23]. Consequently, the vast number of parameters in large AI models can be effectively reduced by employing quantum or hybrid quantum-classical models.
@@
Driven by these advantages, quantum machine learning (QML) has become an active area of research [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Although substantial progress has been made in the development of QML algorithms, there are still few real-world applications that leverage real quantum computers [35, 36, 37, 38, 39]. This gap largely results from the quantum noise issue, which makes results unreliable when many qubits are used [40, 41]. Addressing this noise challenge is complex, and achieving a large-scale, noise-free quantum computer is expected to take considerable time. In light of these obstacles, the development of quantum-inspired classical algorithms that address the challenges of classical machine learning is a valuable direction to advance machine learning techniques [42, 43, 44, 45, 46, 47, 48].
To translate the advantages of quantum machine learning into classical methods, a deeper understanding of their mathematical foundations is required. Currently, such understanding is limited, and research often focuses on empirically demonstrating the success of QML across various datasets. For instance, quantum neural networks have been show to succeed with fewer parameters in tasks like recovery rate prediction and medical image clarification; however, the reasons for this efficiency remain unclear [49, 50]. Our work aims to clarify the mathematical principles of QNNs and explain the origin of these quantum advantages. We will then leverage this understanding to develop a novel classical neural network, achieving high performance with fewer parameters.
@@
We first analyze the mathematical foundations of quantum neural networks with angle encoding and find that using a quantum circuit with angle encoding is mathematically equivalent to constructing a polynomial function, where each term is expressed as a continuous product of the cosine series [47]. In contrast, a quantum circuit with amplitude encoding generates a weight-constrained matrix, whose weight element is again expressed as a summation of the continuous product of the cosine series. The number of trainable variables in this weight-constrained matrix is much smaller than the dimensionality $N$ of the matrix or the input space because the number of quantum gates is typically scaled as $\mathrm{log}_{2}N$ not $N$. Therefore, a quantum circuit with amplitude encoding requires fewer variables than the classical network, whose variable number is typically scaled as $N$. Motivated by this discovery, we develop a classical weight-constrained neural network, which can reduce the number of variables by 100 times compared to the standard neural network (NN). We validate this weight-constrained method by examining its performance on four datasets: MNIST, Fashion MNIST (FMNIST), CIFAR, and TRAFFIC datasets. Our results demonstrate that the weight-constrained method maintains the accuracy of the classical neural network while substantially reducing the number of variables. This approach holds promise for scaling down variable numbers in large AI models such as GPT and BERT. Furthermore, we develop a novel method to enhance the adversarial robustness of the quantum neural network and the classical weight-constrained neural network models by randomly dropping parameterized angles during prediction. We demonstrate that such a method can significantly increase robustness against adversarial attacks.
This paper is organized as follows. Section II presents the mathematical foundations of quantum neural networks with angle and amplitude encodings. Section III develops the theoretical framework of the proposed weight-constrained approach. Section IV demonstrates the effectiveness of applying this method to classical neural networks. Section V compares the performance of weight-constrained neural networks with that of networks employing low-rank matrices. Section VI examines the improved adversarial robustness achieved through our proposed dropout technique. Section VII analyzes the key benefits and broader implications of the weight-constrained methodology. Finally, Section VIII concludes the paper with a summary of findings and potential directions for future research.
Figure 1: Hybrid quantum-classical neural network. (a) A generic quantum circuit with angle encoding. (b) A generic quantum circuit with amplitude encoding.
II Quantum neural networks
@@
To develop quantum-inspired classical neural networks, we explore the underlying mathematics of quantum neural networks with angle and amplitude encodings [1]. After building on this mathematical foundation, we develop a quantum-inspired classical neural network, which utilizes 100 times fewer variables than a standard classical neural network.
@@
We start to analyze the quantum neural network with angle encoding, which adopts the input value as rotation angles in the quantum gates, as shown in Fig. 1(a). The nonlinear relationship between the output of a quantum circuit and the input data $\bf z$} is constructed by applying different rotation gates and CNOT gates in this circuit. Without loss of generality, we consider a quantum circuit shown in Fig. 1(a), which has one data-upload layer $\tilde{R}_{0}$, $M$ rotation layers, and $M$ CNOT layers. The rotation operator $R_{\gamma}$ in Fig. 1(a) is chosen from the collection of $\{R_{x},R_{y},R_{z}\}$. Then, the final quantum state of this quantum circuit is
@@
where $\tilde{R}_{i}$ denotes a matrix constructed from the tensor product of different single-gate rotation matrices within one layer, and $\tilde{C}_{i}$ denotes a matrix constructed from the tensor product of two-gate CNOT matrices. Since $\tilde{R}_{i}$ is composed solely of rotation operators, its matrix elements are either trigonometric functions or zero, whereas the matrix elements of $\tilde{C}_{i}$ are either 0 or 1. Defining
$U=\tilde{R}_{M}\tilde{C}_{M}\cdots\tilde{R}_{1}\tilde{C}_{1}\tilde{R}_{0},$ (2)
then the final state can be expressed as
$|\psi\rangle=U|0000\rangle=\sum_{i}U_{i,1}|b_{i}\rangle,$ (3)
where $b_{i}$ denotes the computational basis states (0000, 0001, 0011, …), and $U_{i,1}$ is the $i$-th element of the first column of $U$, which is either a sum of products of trigonometric functions or zero due to the matrix multiplication. In general, $U_{i,j}$ can be written as
$U_{i,j}=\sum_{s}p_{s}\prod_{i\in\mathbb{N}_{s}}\mathrm{cos}(\theta_{i}/2+q_{i})\prod_{n\in\mathbb{M}_{s}}\mathrm{cos}(z_{n}/2+q_{n}),$ (4)
where $p_{s}$ belongs to {0,1,1,i,i0,1,-1,i,-i}. The symbol $q_{i}$ denotes 0 or $\pi/2$. $\mathbb{N}_{s}$ and $\mathbb{M}_{s}$ are collections of integer numbers.
For a measurement operator $A$ composed of Pauli operators, the circuit output is
$O$ $=$ $\langle\psi|A|\psi\rangle=\sum_{i,j}U^{*}_{i,1}A_{i,j}U_{j,1}=\sum_{s}f_{s}({\bf\theta},{\bf z}),$ (5)
$f_{s}({\bf\theta},{\bf z})$ $=$ $c_{s}\prod_{i\in\mathbb{N}_{1,s}}\mathrm{cos}(\theta_{i}/2+q_{i})\prod_{j\in\mathbb{N}_{2,s}}\mathrm{cos}(\theta_{j}/2+q_{j})\times$
$\prod_{n\in\mathbb{M}_{1,s}}\mathrm{cos}(z_{n}/2+q_{n})\prod_{m\in\mathbb{M}_{2,s}}\mathrm{cos}(z_{m}/2+q_{m}),$
where $c_{s}$ belongs to {1,11,-1}. It is straightforward to find that the output $O$ of a quantum circuit is a polynomial function. Equation (5) shows that the quantum neural network with angle encoding generates a complicated activation function, which could select features better than the classical activation functions, such as ReLU and Tanh. We refer the reader to our previous work [47] for a deeper understanding of this quantum activation function.
Quantum neural network with amplitude encoding is a weight-constrained neural network
@@
where $|b_{i}\rangle$ denotes a pure quantum state, i.e., $|0000\rangle$ and $|0001\rangle$. $|{\bf z}|$ is the norm value of the input vector $\bf z$}. This initial quantum state can be transformed to other states by applying rotation and CNOT gates. For simplicity, we can use an operator $U$ to represent these operations as shown in Fig. 1(b). Then, the final quantum state is written as
@@
Equation (10) shows that the quantum circuit generates a correlation matrix between any two elements in the input data. For example, $W_{1,2}$ represents the correlation between the first and second input features, and $W_{1,100}$ denotes the correlation between the first and the one hundredth input features. If all elements in $W$ are nonzero, the construction of features incorporates long-range and short-range correlations[21].
The weight matrix element $W_{ij}$ is mathematically equivalent to the matrix $U$ defined in Eq. (4). This weight element is given by
$W_{i,j}=\sum_{s}p_{s}\prod_{k\in\mathbb{N}_{s}}\mathrm{cos}(\theta_{k}/2+q_{k}),$ (11)
where $\theta_{k}$ is a rotation angle, $p_{s}\in\{0,1,-1,i,-i\}$. $q_{k}\in\{0,\pi/2\}$, and $\mathbb{N}_{s}$ is a set of integers.
Equation (11) shows that each weight element is a polynomial function, where each term is a continuous product of the cosine series. Therefore, the amplitude encoded quantum neural network is a weight-constrained neural network. In a typical quantum circuit, the number of rotation angles is significantly smaller than the dimensionality of the Fock space it encodes. Consequently, an amplitude-encoded quantum neural network operates with fewer trainable variables than a standard neural network of equivalent representational capacity.
III Quantum-inspired classical network: weight-constrained neural network
@@
Motivated by the discovery that the amplitude encoded quantum neural network is a weight-constrained neural network, we develop a classical counterpart of the weight-constrained neural network, in which the weight variables are constructed as follows. To generate $K$ different weights, we first define $N$ different variables, denoted as $\theta$. Then a combination process is performed, where $r$ variables are chosen from these $N$ variables, resulting in $C(N,r)$ possible combinations. Labeling the $k$-th combination as ${\bf\theta}^{(k)}$, the $k$-th weight $w_{k}$ is defined as
@@
A critical question when adopting the weight-constrained method is how it impacts the performance of a neural network. To answer this question, we examine two distinct two-layer neural networks. In the first classical network, 100 input features, $x$, are connected to an output neuron, $y$, as follows:
@@
where $w_{i}$ represents the weight, sampled from the uniform distribution $\mathcal{U}(-1,1)$. Here, $f(z)$ denotes a standard classical activation function, such as Tanh, Sigmoid, or ReLU. The second classical neural network has the same architecture as the first but employs the weight-constrained method to determine the weights. Since the variance of the constrained weights, as defined in Eq. (12), scales as $2^{-r/2}$, we adjust them to remove the impact of the variance. We redefine the weights as:
$w^{\prime}_{k}=\sqrt{\frac{2^{r}}{3}}\prod_{i\in\mathrm{even},i\leq r}\mathrm{cos}(\theta_{i}^{(k)})\prod_{j\in\mathrm{odd},j\leq r}\mathrm{sin}(\theta_{j}^{(k)}).$ (17)
Then the variance of $w^{\prime}_{k}$ is $\sqrt{1/3}$, matching the variance of the weights $w_{i}$ drawn from $\mathcal{U}(-1,1)$.
Figure 2: Panel (a) and panel (b) plot the distribution $p(z)$ of $z$, defined as $z=\sum_{i}w_{i}x_{i}$, where $w$ denotes the weight, and $x$ denotes the input features. (c) The inverse of Kullback-Leibler divergence $D_{\text{KL}}^{-1}$ of the distribution $p(z)$ as a function of NN at $r=4$, where $N$ denotes the number of variables. (d) $D_{\text{KL}}^{-1}$ of the distribution $p(z)$ as a function of $r$ at $N=16$.
@@
Figure 3: Training results of different neural networks. (a) A summary of the accuracy (ACC) of these five models for MNIST and FMNIST datasets. (b) The architecture of the hybrid quantum-classical neuron network (HNN) and the training results on the MNIST and Fashion MNIST (FMNIST) datasets. (c) The fully connected neuron network (FNN) architecture and the training results on the MNIST and FMNIST datasets. (d) The weight-constrained FNN architecture and the training results on the MNIST and FMNIST datasets. (e) The convolutional neural network (CNN) architecture and the training results on the MNIST and FMNIST datasets. (f) The architecture of the weight-constrained CNN and the training results on the MNIST and FMNIST datasets.
Figures 2(a) and 2(b) display the distribution $p(z)$ of $z$ for both classical neural networks. The input features $x$ are sampled from a uniform distribution $\mathcal{U}(-1,1)$. The black curve represents the distribution for the classical neural network with weights sampled from $w\in\mathcal{U}(-1,1)$. The other curves show results for the weight-constrained network across different pairs of $(N,r)$ used in constructing the combination sets. As shown in Figs. 2(a) and 2(b), both networks exhibit dome-shaped distributions centered at $z=0$. Including a uniformly distributed bias term does not affect the central position of these distributions.
Although the mean and variance of the distributions for the weights from the uniform distribution and the weight-constrained method are the same, the weight-constrained network displays a narrower distribution width and a more pronounced peak compared to the network using uniformly distributed weights. Moreover, we observe that the weight distribution of the weight-constrained network shows a strong dependence on $r$ but a weak dependence of $N$. For instance, at $r=4$, increasing the value of $N$ from 14 to 20 only slightly raises $p(z=0)$ from 0.06 to 0.08. In contrast, at $N=16$, increasing $r$ from 4 to 10 causes a significant enhancement of $p(z=0)$. This result implies that the performance of the weight-constrained network has a strong dependence on $r$ but a weak dependence of $N$.
To quantify the performance of a network, we evaluate its expressibility, which characterizes the ability of a network to represent or predict arbitrary real value. In the ideal case, the variable $z$ can take any real value within the range $(-\infty,\infty)$, and its probability distribution $p(z)$ is uniform. Such a case corresponds to the strongest expressibility. However, when constraints are imposed on both the weights $w$ and inputs $x$, the resulting distribution $p(z)$ becomes dependent on $z$, thereby reducing the expressibility. To quantify this deviation, we employ the Kullback-Leibler (KL) divergence, $D_{\mathrm{KL}}$, which measures the difference between $p(z)$ of the network and $q(z)$ of the ideal case:
$D_{\text{KL}}=\int_{-\infty}^{\infty}dzp(z)\log\left[\frac{p(z)}{q(z)}\right].$ (18)
In practical computations, we should restrict the integration domain to (R,R)(-R,R) with a large cutoff $R$, for which $q(z)=1/(2R)$. Under this approximation, the KL divergence becomes
$D_{\text{KL}}=\int_{-R}^{R}dzp(z)\log p(z)+\log(2R),$ (19)
when $p(z)$ closely resembles the uniform distribution of the ideal case, $D_{\text{KL}}\approx 0$; conversely, when $p(z)$ collapses to a delta function $\delta(z)$, we obtain $D_{\text{KL}}=\log(2R)$. Therefore, $D_{\text{KL}}$ serves as an inverse measure of expressibility — the smaller the KL divergence, the greater the expressibility of a network.
In our calculations, we use the output distribution $p(z)$ of a network with weights sampled from U(1,1)U(-1,1) as the reference distribution $q(z)$. This choice is justified because this unrestricted network possesses higher expressibility than the weight-constrained network (see Appendix F for details). Figures 2(c) and 2(d) plot the inverse KL divergence, $1/D_{\text{KL}}$ (blue circles), as a function of $N$ (at $r=4$) and $r$ (at $N=16$), respectively. The results indicate that $1/D_{\text{KL}}$ decreases slightly with increasing $N$ for $N<18$ and becomes independent of $N$ for $N>18$. In contrast, $1/D_{\text{KL}}$ decreases rapidly as $r$ increases. This behavior suggests that the expressibility of the weight-constrained network depends weakly on $N$, while it increases significantly as $r$ decreases.
IV Practicality of the weight-constrained method